OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 10x - 6, find f(g(x)). How do I solve that?! Step by step instruction, please; don't just do it for me. I really need to know HOW it's done.

OpenStudy (anonymous):

Hint: (-4x+7(10x-6))

OpenStudy (anonymous):

How is that @thetechnomage

OpenStudy (anonymous):

My options are: a. -40x + 70 b. -40x + 64 c. -40x + 24 d. -40x + 31 So ... yeah :/ Still confused.

OpenStudy (anonymous):

Sorry my bad, it has to be -40x + 31

OpenStudy (anonymous):

I did mistake in last step

OpenStudy (anonymous):

Okay! Thanks so much!

OpenStudy (anonymous):

put the function g(x) into the function f(x) which means that you put 10x-6 instead of the x in f(x) So, to do that: f(g(x)) = -4 (10x-6) + 7 = -40x +24 +7 = -40x + 31

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

That makes SO much more sense to me! Thank you so much!

OpenStudy (anonymous):

np